What is the sum of the first 75 consecutive odd numbers?
Answer by Alan3354(69443)

What is the sum of the first 75 consecutive odd numbers?
----------------
The 1st 76 odd numbers are from 1 to 151
Make 38 pairs, 1 & 151, 3 & 149, etc
Each pair is 152
38*152 = 5776 (= 76^2)
Subtract 151
= 5625 (= 75^2)